a pharmacist wishes to mix a solution that is 4% of minoxidil. She has on hand 90 ml of 2% solution and she wishes to add some 6% solution to obtain the desired 4% solution. How much 6% should she add?
Answer by stanbon(75887)

a pharmacist wishes to mix a solution that is 4% of minoxidil. She has on hand 90 ml of 2% solution and she wishes to add some 6% solution to obtain the desired 4% solution. How much 6% should she add?
-------
Equation:
mino + mino = mino
0.02*90 + 0.06x = 0.04(90+x)
-------
Multiply thru by 100 to get:
2*90 + 6x = 4*90 + 4x
2x = 2*90
x = 90 ml (amt. of 6% solution to add)
==================
Cheers,
Stan H.